OpenStudy (anonymous):

If a is uniformly distributed over [−28,28], what is the probability that the roots of the equation x2+ax+a+80=0

OpenStudy (anonymous):

what is the probability that the roots of the equation what?

OpenStudy (anonymous):

what it the probability that the roots of the equation are real?

OpenStudy (anonymous):

are both real?

OpenStudy (anonymous):

oh ok we can solve that using the quadratic formula

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what would be the b value of the equation. is it a(x+1)

OpenStudy (anonymous):

\[x^2+ax+a+80=0 \] \[x=\frac{-a\pm\sqrt{a^2-4a-320}}{2}\]

OpenStudy (anonymous):

no if you use the quadratic formula, \(a=1,b=a, c=a+80\) different \(a\) of course

OpenStudy (anonymous):

if these are to be real numbers, that means you must have \[a^2-4a-300\geq 0\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

by some miracle this factors as \[(a-20)(a+16)\geq 0\] so we can actually solve

OpenStudy (anonymous):

how did i get \(a^2-4a-320\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or how did i get the whole equation?

OpenStudy (phi):

how did you get that? the stuff under the square root has to be 0 or positive, else you get imaginary numbers when you take the root.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok quadratic formula tells you the solution to \(ax^2+bx+c=0\) is \(x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) in your case \(a=1,b=a,c=a+80\)

OpenStudy (anonymous):

when you compute you get \[x=\frac{-a\pm\sqrt{a^2-4a-320}}{2}\] in your case so for these to be real, the discriminant \(a^2-4a-320\) must be greater than or equal to zero, otherwise you have a negative number under the radical

OpenStudy (anonymous):

oh, what @phi said

OpenStudy (anonymous):

okay so after i find my values what do I do?

OpenStudy (anonymous):

your last job is to solve for \(a\) \[a^2-4a-320\geq 0\]

OpenStudy (anonymous):

i found my values and they are a=20 and a+-16

OpenStudy (anonymous):

this factors as \((a-20)(a+16)\geq 0\)

OpenStudy (anonymous):

a=-16

OpenStudy (anonymous):

hold on, you are not solving \((a-20)(a+16)=0\) but rather \[(a-20)(a+16)\geq 0\]

OpenStudy (anonymous):

the zeros are \(-16\) and \(20\) for sure but you want to know the interval over which it is positive since this is a quadratic with leading coeffient positive, it will be positive outside the zeros, in other words if \(a\leq -16\) or \(a\geq 20\)

OpenStudy (anonymous):

your final job is to find what portion of the interval \([-28,28]\) satisfies \(a\leq -16\) or \(a\geq 20\)

OpenStudy (anonymous):

okay so i would have to find the interval of the it by integrating it?

OpenStudy (anonymous):

not necessary, just look

OpenStudy (anonymous):

favorable part is \([-28,-16]\cup [20,28]\)

OpenStudy (anonymous):

you can eyeball the total length

OpenStudy (anonymous):

12 and 8

OpenStudy (anonymous):

right, for a total of 20

OpenStudy (anonymous):

thanks just found the answer. thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hope it was \(\frac{20}{56}\)

OpenStudy (anonymous):

yea